# results on dynamics of bungee set of composite entire functions in the eremenko-lyubich class

### v3

@@ Abstract.
In this paper, we have discussed the dynamics of composite entire functions in terms of relationship between bungee set, repelling periodic points (to be denoted by $RP$) and rationally indifferent fixed point set. We have established relation between the dynamics of composite function and the functions taken for composition. We showed that the union of their bungee sets contains bungee set of the composite function. We also showed that $RP$ set of composite functions contains the $RP$ set of functions used for the composition. We have mostly dealt with the permutable functions of class $\mathcal{B}$. These results hold true when we have a composite function in which one function is a periodic translation of another.
@@ 1. Introduction
When we iterate a holomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n-$times, we observe that the points of the complex plane move differently, which we call the dynamics of the point where we denote $f^{n}$ to be $n-$th iteration of a function, where $n\in\mathbb{N}$. As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour). This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
@@
There is yet another classification of these points based on the nature of their orbits; viz., the filled- in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded). These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$. In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$. $K(f)$ comprises of Fatou components in which orbits under iteration stays bounded, along with repelling periodic points and rationally indifferent periodic points that lies inside Julia set $J(f).$ It is to be noted that the bungee set and escaping set must have a finite intersection with $J(f)$ (see [5, 16]). The escaping set too has a finite intersection with the Fatou set when the Fatou component is a Baker domain. Knowing all these sets for an entire function, we were curious about what happens when we compose two functions and how these sets relate to the composition with those of individual functions. We tried to see what happens to the dynamics of transcendental functions under compositions as composition of two transcendental functions is again transcendental.
@@
Simplifying the situation further, we have taken transcendental functions $f,g$ of Eremenko Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded). Additionally, this class does not have any intersection of the Fatou set with the bungee set, in other words there is no wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ Since our aim was to establish a relationship between different sets of functions and their composition and it would be quite difficult to establish any relation where there is no algebraic relation among these functions, we have focused on commuting(permutable) functions, that is, $f\circ g=g\circ f$, and $g=f+C$ where $C$ is the period of $f$. The following results are proven in this paper:
Let $f,g\in\mathcal{B}$ be commuting functions. The bungee set of composed functions is contained in the union of the bungee set of functions used for composition. The set of repelling periodic and indifferent fixed points of the composed function $RP(f\circ g)$ contains the intersection of repelling periodic and indifferent fixed points of these functions; that is, $RP(f\circ g)\supset RP(f)\cap RP(g)$.
@@
We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [18]. Also, $\mathcal{S}\subset\mathcal{B}$. $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes under composition with $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$ [2]; therefore, in this case, the result clearly holds.
@@ 1.1. Observations
While developing the relationships among the functions and the individual functions, we also found that the bungee set and escaping set exhibit similar characteristics for functions of class $\mathcal{B}$:
@@ 2. Main results
To begin with the commuting entire functions, we first understand the nature of these sets when functions have no finite asymptotic values in their singular set.
@@
If $f$ and $g$ are commuting functions having no finite asymptotic values, then $I(f)$ is completely invariant under $g$ and vice versa.
@@
Theorem 2.2.
Suppose $f$ and $g$ are commuting entire functions having no finite asymptotic values. Then $K(f)$ and $BU(f)$ are completely invariant under $g$ and vice versa.
Proof.
$g(K(f))$ is, in general forward invariant, because if we assume a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|g(f^{n}(w))|=|f^{n}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$, that proves the result. Now, let $z_{0}\in g^{-1}(K(f)).$ This implies that $g(z_{0})\in K(f)$ and $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ As these functions do not have any finite asymptotic values, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$. Thus, we have that $K(f)$ is completely invariant under $g$ when the functions have no finite asymptotic values. Similarly $K(g)$ is also completely invariant under $f$ as well.
Now bungee set, $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know from preceding argument $K(f)$ is completely invariant under $g$ and from Theorem 2.1 $I(f)$ is completely invariant under $g.$ Hence, $BU(f)$ is also completely invariant under $g$. Similarly, $BU(g)$ will be completely invariant under $f$. ∎
Lemma 2.3.
The filled-in Julia set of the composite function $f\circ g$ cannot be larger than the intersection of filled-in Julia set of the individual functions which are commuting and having no finite asymptotic values.
Proof.
We assume that $K(f\circ g)\supset K(f)\cap K(g)$. Thus, there might be points in $K(f\circ g)$ that have a finite intersection with $I(f)\cup I(g)\cup BU(f)\cup BU(g)\cup A\cup B$ where $A:K(f)\backslash K(f)\cap K(g)$ and $B:K(g)\backslash K(f)\cap K(g)$. If $z_{0}\in I(f)\cup I(g)\cup BU(f)\cup BU(g)$, then iterate of $z_{0}$ under $(f\circ g)^{n}$ cannot be bounded because $f^{n}\circ g^{n}(z_{0})$ as $n\rightarrow\infty$ will either remain oscillatory or will tend to $\infty$ (Theorem 2.2 and Theorem 2.1).
Now, let $w_{0}\in A$ and apply $f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ as $m\rightarrow\infty$, which is bounded because $g(K(f))\subset K(f)$. Now, as $w_{0}\not\in K(g)$, and after interchanging roles of $f$ and $g$ due to commutation, we observe that $g^{m}\circ f^{2m}\circ g^{m}(w_{0})$ cannot be bounded (as $w_{0}$ will either be in $BU(g)$ or $I(g)$ and $f(BU(g)\subset BU(g)$, $f(I(g)\subset I(g)$). Hence, $g^{m}\circ f^{2m}\circ g^{m}(w_{0})=f^{m}\circ g^{2m}\circ f^{m}(w_{0})$ in which RHS is bounded while LHS is not which is a contradiction. Similar arguments may be given when $w_{1}\in B$. Hence, $K(f\circ g)\not\supset K(f)\cap K(g)$. ∎
The next result provides condition for equality of bungee set, filled-in Julia set and escaping set for commuting functions.
Theorem 2.4.
Commuting transcendental entire functions with no finite asymptotic values have equal bungee sets, filled-in Julia sets and escaping sets.
Proof.
We first establish that $BU(f\circ g)\subset BU(f)\cap BU(g).$ Let $z_{0}\in BU(f\circ g)$. By definition, there are two subsequences $\{m_{k}\}$ and $\{n_{k}\}$ both tending to $\infty$ s.t.,
$|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R\ \text{for some}\ R$
$|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty$
So, for $\{m_{k}\}$ subsequence, $|(f\circ g)^{m_{k}}(z_{0})|=|f^{m_{k}}\circ g^{m_{k}}(z_{0})|<R$ for some $R.$ As $f$ does not have any finite asymptotic value, so $|g^{m_{k}}(z_{0})|<R_{1}$ for some $R_{1}$. We observe same thing for $f$ as well i.e., $|f^{m_{k}}(z_{0})|<R_{2}$ for some $R_{2}$ (because $|f^{m_{k}}\circ g^{m_{k}}(z_{0})|=|g^{m_{k}}\circ f^{m_{k}}(z_{0})|<R$ as $f,g$ are commuting).
Now for $\{n_{k}\}$ subsequence, $|(f\circ g)^{n_{k}}(z_{0})|=|f^{n_{k}}\circ g^{n_{k}}(z_{0})|\rightarrow\infty.$ We have 3 possible cases: $|g^{n_{k}}(z_{0})|\rightarrow\infty$ or $|g^{n_{k}}(z_{0})|\rightarrow R_{3}$ for some $R_{3}$ or $g^{n_{k}}(z_{0})$ itself exhibits bungee behaviour. If we have the second case, it means $z_{0}\in K(g)$ (as this will imply all subsequences of $g^{j}(z_{0})$ are bounded where $j$ is some arbitrary index unrelated to $m_{k}$ and $n_{k}$). As a result, $g^{j}\circ f^{2j}\circ g^{j}(z_{0})$ will remain bounded (because $g^{j}\circ f^{2j}(K(g))\subset g^{j}(K(g))$ which is bounded). Hence, this case isn’t possible and $z_{0}\not\in K(g)$. On similar lines, it can be seen that $z_{0}\not\in K(f)$ where similar arguments follow after interchanging roles of $f$ and $g$ due to commutation. Now, when we have either of the first or the third case, it means the same thing. Both these cases essentially implies that $z_{0}\in BU(g)$. From Theorem 2.2, we have $g^{n_{1k}}\circ f^{2n_{1k}}\circ g^{n_{1k}}(z_{0})$ exhibits bungee behaviour which is as expected. Due to commutation, we also conclude that $z_{0}\in BU(f)$ as well. Hence, $BU(f\circ g)\subset BU(f)\cap BU(g)$.
Given the preceding proof, we ask a question:
Question: What can we say about points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g),$ where will it be situated?
Answer: As the functions $f$ and $g$ do not have any finite asymptotic values, therefore, points of the set $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $K(f\circ g).$ Also, $I(f\circ g)\subset I(f)\cup I(g)$ [14] and as $(BU(f)\cup BU(g))\cap(I(f)\cup I(g))=\emptyset,$ it means that $BU(f)\cup BU(g)\setminus BU(f)\cap BU(g)$ cannot intersect with $I(f\circ g)$ either. This implies that the only place it will be contained is $BU(f\circ g)$. However, if it happens, then such points must be contained in $BU(f)\cap BU(g)$ implying that no such points exist. This is only possible if $BU(f)=BU(g)$.
Similar question can be asked for filled-in Julia sets as well, (as it was shown in Lemma (2.3) that $K(f\circ g)\subset K(f)\cap K(g)$). And the answer is: Let $w_{0}\in K(f)\cup K(g)\ \backslash\ K(f)\cap K(g)$. These points won’t intersect with $I(f\circ g)$ (as $I(f\circ g)\subset I(f)\cup I(g)$ and $K(f)\cup K(g)\cap I(f)\cup I(g)=\emptyset$ [14]) nor with $BU(f\circ g)$ (as $BU(f\circ g)=BU(f)=BU(g)$ and $BU(f\circ g)\cap K(f)\cup K(g)=\emptyset)$ So, they will intersect with $K(f\circ g)$ which means $w_{0}\in K(f)\cap K(g)$ which is a contradiction. Hence, we conclude no such points are possible. Therefore, $K(f\circ g)=K(f)=K(g)$ when there are no finite asymptotic values for these commuting functions.
Hence, given the following results:
$K(f\circ g)$ $=K(f)=K(g)$
and
(2.1) $BU(f\circ g)$ $=BU(f)=BU(g)$
one can conclude that $I(f\circ g)=I(f)=I(g)$, for we know that $K(f)\cup I(f)\cup BU(f)=\mathbb{C}=K(f\circ g)\cup I(f\circ g)\cup BU(f\circ g)=K(g)\cup I(g)\cup BU(g)$ and from the previous results (2) we can conclude that $I(f\circ g)=I(f)=I(g)$. ∎
@@
Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ do not have any finite asymptotic values.
One may also consider the following functions for illustration of above result:
$f_{1}(z)=1+z+e^{z}$ which would be commuting with $g_{1}(z)=1+z+e^{z}+2\pi i$; $f_{2}(z)=z+sin(z)$ which would be commuting with $g_{2}(z)=z+sin(z)+2\pi.$

### v4

@@ Abstract.
In this paper, we have discussed the dynamics of composite entire functions in terms of relationship between bungee set, escaping set and filled-in Julia set. We have established some relation between the dynamics of composition of entire functions and the functions taken for composition. We have shown that the union of the bungee set of two entire functions contains the bungee set of the composite function. In addition, it is shown that the filled-in Julia set of composite entire functions contains the filled-in Julia set of functions used for the composition. The results have been illustrated with several examples. We have mostly dealt with permutable(commuting) functions.
@@ 1. Introduction
When we iterate a holomorphic function $f:\mathbb{C}\rightarrow\mathbb{C}$, $n$-times, we observe that the points of the complex plane move differently, which we call the dynamics of the point. Here, $f^{n}$ denotes the $n$-th iteration of function $f$, where $n\in\mathbb{N}$. As, not all points have same dynamics, we therefore have different classifications of points of the complex plane $\mathbb{C}$; one classification is based on the nature of neighbourhood of a point $z_{0}$ (say); they are said to be in the Fatou set $F(f)$ (set of points whose neighbourhood exhibits stable behaviour) and Julia set $J(f)$ (set of points whose neighbourhood exhibits chaotic behaviour). This was broadly the classification we generally see when we talk about dynamics of points in complex plane.
@@
There is yet another classification of these points based on the nature of their orbits; viz., the filled-in Julia set $K(f)$ (set of points whose orbits are bounded), escaping set $I(f)$ (set of points whose orbits tend to infinity), and bungee set $BU(f)$ (the orbits of such points neither remain bounded nor tend to infinity; only there exists a subsequence that remains bounded and another one that is unbounded). These sets are disjoint among themselves and their union $BU(f)\cup I(f)\cup K(f)=\mathbb{C}$. In [16], it was shown that if a Fatou component intersects with $BU(f)$ then that component becomes a wandering domain contained in $BU(f)$. It is to be noted that the bungee set and escaping set must have a non empty intersection with $J(f)$ (see [5, 16]). The escaping set too has a non empty intersection with the Fatou set when the Fatou component is a Baker domain [14], or when the Fatou set has a multiply connected wandering domain [19]. Knowing all these sets for an entire function, we were curious about what happens when we compose two such functions and how these sets relate to the composition with those of individual functions. As composition of two transcendental functions is again transcendental, we have tried to analyse what happens to the dynamics of transcendental functions under compositions.
@@
Simplifying the situation further, we have taken transcendental functions $f,g$ in the Eremenko-Lyubich class $\mathcal{B}$ which means that the functions of this class have a bounded set of singular values (i.e., the set of all asymptotic values and critical values and their finite limit points is bounded). Recall the notion of an oscillatory wandering domain.
Definition 1.1.
Given a transcendental entire function $f$, a domain $U\subset F(f)$ is called an oscillatory wandering domain of $f$, if $U\subseteq BU(f)$, that is, $\{\infty,d\}$ is a limit function of $\{f^{n}\}$ on $U$ for some $d\in\mathbb{C}$, [17].
Additionally, class $\mathcal{B}$ may have an intersection of the Fatou set with the bungee set, in other words there can be oscillating wandering domain which also lies inside the bungee set for $f\in\mathcal{B}.$ As our aim was to establish a relationship between different sets of functions and their compositions, it would be very difficult to establish any such relation in the absence of some algebraic relation among these functions. Hence, we have focused on permutable (commuting) functions, that is, functions $f,g$ satisfying $f\circ g=g\circ f.$ Also, we have considered entire functions $f$ and $g$ related by $g=f+C$ where $C$ is the period of $f$. The following results are proven in this paper:
@@
Let $f,g\in\mathcal{B}$ be commuting functions. The bungee set of composed function is contained in the union of the bungee set of functions used for composition. We know some results for exponential functions which belong to Speiser class $\mathcal{S}$ (such class of functions has finite number of elements in their singular set) [20]. Also, $\mathcal{S}\subset\mathcal{B}$. It was shown in [2] that for $f=e^{\lambda z},\lambda\neq 0$, a function $g$ that commutes with $f$ is some finite iteration of $f$ itself, that is, $g=f^{n}$ for some $n\in\mathbb{N}$. Hence, in this case, the result clearly holds, in fact, we obtain $BU(f)=BU(g)=BU(f\circ g).$
@@ 1.1. Observations
While developing the relationships among the sets of composite functions and the individual functions, we also found that the bungee set and escaping set exhibit similar characteristics for functions of class $\mathcal{B}$:
@@ 2. Main results
Property A : $f$ (respectively $g$) has the property that it sends orbits escaping under $g$ (respectively, $f$) to $\infty.$
To begin with the commuting entire functions, we first understand the nature of these sets when functions satisfies Property A.
@@
If $f$ and $g$ are commuting functions satisfying Property A, then $I(f)$ is completely invariant under $g$ and vice versa.
@@
Suppose $z_{0}\in I(f).$ Then, $|f^{n}(z_{0})|\rightarrow\infty$. Now $|g(f^{n}(z_{0})|$ can either tend to $\infty$ or be bounded. Using Property A, $|g(f^{n}(z_{0})|\rightarrow\infty$. As $f$ and $g$ are commuting so, $|g(f^{n}(z_{0}))|$ = $|f^{n}(g(z_{0}))|\rightarrow\infty.$ This implies that $I(f)$ is forward invariant under $g$. It is easy to see that $g^{-1}(I(f))\subset I(f)$. Hence, $I(f)$ is completely invariant under $g$. As $f$ and $g$ are commuting, their roles can be interchanged and hence, $I(g)$ is completely invariant under $f$ as well. ∎
Remark 2.2.
It follows from above result that $I(f)$ and $I(g)$ are completely invariant under $f\circ g$ as well.
Theorem 2.3.
Suppose $f$ and $g$ are commuting entire functions satisfying Property A. Then $K(f)$ and $BU(f)$ are completely invariant under $g$ and vice versa.
Proof.
$K(f)$ is in general forward invariant under $g,$ because if we take a point $w\in K(f)$, then we have $|f^{n}(w)|<R$ for some $R.$ It is clear that $|g(f^{n}(w))|=|f^{n}(g(w))|<R_{1}$ for some $R_{1}$ which implies that $g(w)\in K(f)$ that proves the result. Now, let $z_{0}\in g^{-1}(K(f)).$ This implies that $|f^{n}(g(z_{0}))|<L$ for some $L$ and because of commutation, $|g(f^{n}(z_{0}))|<L.$ Using Property A, $|f^{n}(z_{0})|<L_{1}$ for some $L_{1}$, which means $z_{0}\in K(f)$. Thus, we obtain that $K(f)$ is completely invariant under $g$. Similarly, $K(g)$ is also completely invariant under $f$ as well. ∎
Theorem 2.4.
Suppose $f$ and $g$ are commuting entire functions satisfying Property A. Then $BU(f)$ is completely invariant under $g$ and vice versa.
Proof.
Using the dynamical partition of the complex plane, we obtain $BU(f)=\mathbb{C}\backslash(I(f))\cup K(f)).$ We know from preceding argument that $K(f)$ is completely invariant under $g$ and from Theorem 2.1, $I(f)$ is completely invariant under $g.$ Hence, on combining these results, we conclude that $BU(f)$ is also completely invariant under $g$. On similar lines, $BU(g)$ will also be completely invariant under $f$. ∎
Remark 2.5.
It follows from above result that $I(f),$ $I(g);K(f),K(g);$ and $BU(f),BU(g)$ are completely invariant under $f\circ g$ as well.
The next result provides a relation between the filled-in Julia set of composite entire functions and the filled-in Julia set of the individual functions.
Lemma 2.6.
The filled-in Julia set of the composite function $f\circ g$ contains the intersection of the filled-in Julia set of the individual functions which are commuting and satisfying Property A.
Proof.
Suppose that $z_{0}\in K(f)\cap K(g)$. Consider the array of following sequences:
$\begin{array}[]{ccccc}f(g(z_{0}))&f^{2}(g(z_{0}))&f^{3}(g(z_{0}))\cdots\\ f(g^{2}(z_{0}))&f^{2}(g^{2}(z_{0}))&f^{3}(g^{2}(z_{0}))\cdots\\ \vdots&\vdots&\vdots\\ f(g^{k}(z_{0}))&f^{2}(g^{k}(z_{0}))&f^{3}(g^{k}(z_{0}))\cdots\\$ $\vdots$$\vdots&\vdots\par\end{array}$
Now, the first row of the array will converge to say, $l$ for, if it does not, then $g(z_{0})\in I(f)$ which implies that $z_{0}\in I(f)$, a contradiction. Using continuity of $g$, we can say that second row will converge to $g(l)$ and so on. Let $X$ be the set containing all these limit points. The following are the possible cases for set $X$.
Case 1: $l$ is periodic point of $g$. In particular, suppose that $l$ is a fixed point of $g$. Hence, every sequence of the array will be bounded which shows that $z_{0}\in K(f\circ g)$.
Case 2: $X$ is infinite and bounded. Then, by Bolzano-Weierstrass theorem, $X$ has a limit point which also shows that $z_{0}\in K(f\circ g)$.
Case 3: $X$ is unbounded, so that there exists a subsequence which escapes to infinity. This in turn implies that $z_{0}\in I(f\circ g)\subset I(f)\cup I(g)$ (see [14]) which leads to a contradiction.
From all the above cases, we obtain that $K(f)\cap K(g)\subset K(f\circ g).$
The next result connects bungee set of composite entire function with that of individual functions.
Theorem 2.7.
Suppose $f$ and $g$ are commuting transcendental entire functions satisfying Property A. Then, $BU(f\circ g)\subset BU(f)\cap BU(g).$
Proof.
Suppose $z_{0}\in BU(f\circ g).$ Then, there exists two subsequences $\{n_{k}\},\{m_{k}\}$ and a constant $M>0$ such that $(f\circ g)^{n_{k}}(z_{0})\to\infty$ as $k\to\infty$ and $|(f\circ g)^{m_{k}}(z_{0})|<M\mbox{ for all }k\in\mathbb{N}$. It is enough to show that $z_{0}\in BU(g)$ (as the proof of $z_{0}\in BU(f)$ follows on similar lines). In order to show this, we prove that $z_{0}$ cannot be in $I(g)$ as well as $K(g).$ We divide the proof into several cases:
Case(i): Suppose that $g^{n_{k}}(z_{0})\to\infty$ and $g^{m_{k}}(z_{0})\to\infty$ as $k\to\infty$.
We can assume that $g^{n}(z_{0})\to\infty$ as $n\to\infty$ (for, if there exists a subsequence $\{l_{k}\}$ such that $g^{l_{k}}(z_{0})$ stays bounded as $k\to\infty$ then this will clearly imply that $z_{0}\in BU(g)$). Also, using Theorem 2.1, $f^{k}(I(g)\subset I(g)$ for every $k\in\mathbb{N}$. In particular, $f^{m_{k}}(z_{0})\in I(g)$ which implies that $(g\circ f)^{m_{k}}(z_{0})\to\infty$ as $k\to\infty$ which leads to a contradiction.
Case(ii): Suppose that $g^{n_{k}}(z_{0})\to b$ and $g^{m_{k}}(z_{0})\to\infty$ as $k\to\infty$. This again shows that $z_{0}\in BU(g)$ and the result follows.
Case(3): Suppose that $g^{n_{k}}(z_{0})\to\alpha$ and $g^{m_{k}}(z_{0})\to\beta$ as $k\to\infty.$ It follows that $z_{0}\in K(g)$ (for, if there exists a subsequence $\{l_{k}\}$ for which $g^{l_{k}}(z_{0})\to\infty$ as $k\to\infty$ then we are back to Case(2) and hence $z_{0}\in BU(g)$). Using Theorem 2.1, $f(K(g)\subset K(g)$. In particular, $f^{n_{k}}(z_{0})\in K(g)$, i.e., there exists $L>0$ such that $|g^{n}(f^{n_{k}}(z_{0}))|<L\mbox{ for all }n\in\mathbb{N}$. This implies that $|(f\circ g)^{n_{k}}(z_{0})|<L\mbox{ for all }k\in\mathbb{N}$ which is again a contradiction. Thus, from the above cases, we conclude that $z_{0}\in BU(g).$ The proof of $z_{0}\in BU(f)$ follows on similar lines. Hence, we obtain that $BU(f\circ g)\subset BU(f)\cap BU(g)\subset BU(f)\cup BU(g)$. ∎
Given the preceding proof, we ask a question:
Question: What can we say about points of the set $K(f)\cup K(g)\setminus K(f)\cap K(g),$ where will it be situated?
@@
Consider $f(z)=1+z+e^{-z}$(Fatou map) and $g(z)=1+z+e^{-z}+2\pi i.$ It can be easily checked that $f$ and $g$ satisfies Property A.
One may also consider the following functions for illustration of above result. They also satisfy Property A.
$f_{1}(z)=1+z+e^{z}$ which would be commuting with $g_{1}(z)=1+z+e^{z}+2\pi i$; $f_{2}(z)=z+\sin(z)$ which would be commuting with $g_{2}(z)=z+\sin(z)+2\pi.$